# Towards an Accessible, Noninvasive Micronutrient Status Assessment Method: A Comprehensive Review of Existing Techniques

### v1

@@ 3 Biofluid Analysis Methods for Micronutrient Status Assessment
This section describes how biofluid analysis has been leveraged to assess micronutrient status in humans. These assessment methods target particular biomarkers within a biofluid that are indicators of micronutrient status. Reliable biomarkers are a research challenge themselves (which AI techniques may address [36]), but clinical literature shows that micronutrient biomarkers are more established and specific than macronutrient biomarkers [43]. Most micronutrients have one or two specific biomarkers associated with their circulating status that are considered to be the ’gold standard’ for status assessment. Although some gold standards are still debated, their existence makes the evaluation of novel assessment methods more straightforward.
3.1 Clinical Biochemical Analysis
Clinical biochemical analysis involves laboratory testing of biomarkers (biological compounds) found in urine, blood, or other biosamples [142]. Results can be influenced by several factors and need to be interpreted in the context of other factors of the patient’s health. Additionally, biochemical testing is often time and resource-intensive. Despite this, biochemical analysis describes clinical gold-standard methods for quantifying the circulating micronutrient status in the body [189, 69]. These methods can be roughly separated into assays and liquid chromatography (LC)-coupled spectroscopy. Assays work on the principle that specific micronutrients are needed for the growth of certain bacteria, and this growth can be measured to indicate the amount of a micronutrient present in a sample [189]. Assays were previously the widely-accepted gold standard, but improvements in LC and spectroscopy highlight their relatively poor precision and accuracy. These flaws have relegated assay methods to be used mostly during screening or in resource-constrained testing, except for some micronutrients, where they remain the standard. A common assay applied in studies is the enzyme-linked immunosorbent assay (ELISA). An ELISA test is used for measuring antibodies in blood, and is a useful clinical screening tool for further testing [88]. While not the gold-standard for the assessment of micronutrient status, ELISA tests can yield valuable results for studies that are time or resource limited.
Modern gold-standards overwhelmingly apply LC-coupled spectroscopy [189, 69] (Tables 17 to 19). LC is defined as "a separation process used to isolate the individual components in a mixture" [29]. High-performance LC (HPLC) uses pressure to facilitate the separation process, reducing the time required. It is commonly coupled with mass spectroscopy (MS) in gold-standard approaches [189, 69]. Spectroscopy is the "investigation and measurement of spectra produced by matter interacting with or emitting electromagnetic radiation" [2]. Every molecule reacts to the applied radiation in a unique way that allows us to "detect, determine, or quantify the molecular and/or structural composition of a sample" [2]. MS is the most important subfield of spectroscopy to understand for biochemical analysis. It measures the mass-to-charge ratio of the molecules in a sample as a way to determine and quantify the composition of molecules in the sample [2]. This is done by vaporizing the molecules in a sample into gas-phase ions, which are then sorted by their mass-to-charge ratios. We will discuss other forms of spectroscopy and their utility for micronutrient status assessment further in Section 3.4.
Matrices are the biosamples that are the subjects of the aforementioned methods of analysis. Most are blood based, but in a few cases urine is used in the gold-standard (mostly for water-soluble vitamins) [69]. Blood matrices are whole blood, washed red blood cells, plasma, and serum. Whole blood is blood as it is from the vein. Washed red blood cells are red blood cells that have been separated from the other components of blood such as plasma, platelets, and white blood cells [81]. Plasma is obtained by adding an anticoagulant to whole blood and placing it in a centrifuge [165]. Serum is obtained similarly to plasma, except the blood is allowed to clot before centrifuging.
@@ 3.1 Clinical Biochemical Analysis
As we proceed to review the non-clinical methods of biofluids analysis (Sections 3.2 to 3.5), we will be paying attention to what each approach claims to assess, the method by which they conduct this assessment, how their solution was evaluated, and the clinical relevance of their implementation. This last element includes the target biomarker and the concentrations of that biomarker that are evaluated. A clinically-relevant method should closely align with the gold-standard on these factors (Tables 17 to 19). To aid future work that may wish to integrate or innovate on a particular method, we explicitly mention when a study does not demonstrate this agreement. We also provide tables that group together similar works and summarize the pertinent details and quantitative results of their evaluation, if available.
3.2 Assay-Based Technology
@@
One of the largest areas of work applicable to micronutrient status explored by this review is quantitative assays. Although we cover many different types of quantitative assays, a comprehensive review of lateral flow quantitative assays by Urusov et al. [174] provides a robust background for how these devices work. Lateral flow (immunochromatographic) assays indicate that a target compound is either present in the sample or present in excess of a particular threshold, usually via staining on the test membrane. These types of assays are useful for tests that benefit from quick conclusions (e.g. pregnancy tests). The most prevalent approach for extracting quantitative information from lateral flow assays is via optical signal registration [174]. This method involves the analysis of absorbed and reflected light from the test surface and the staining upon it, similar to the practice of spectrophotometry. The test/control ratio is a common metric used to quantify the magnitude of this staining relative to a control or reference area. Some commercially-available devices are limited to automatically confirming the presence of the test line, while others can use line intensity to calculate analyte content. Urusov et al. [174] note that portability has become a recent focus in this market, and this is not just limited to specialized devices in a portable form factor. Smartphones have been successfully used for optical signal registration, even with fluorescent labels that decrease detection limits. Some manufacturers provide their own smartphone apps for quantitative analysis, but controlling for lighting and positioning is a challenge. To address this issue, another approach is the use of a standardized scanning device to collect image data and the off-device analysis of the image by specialized software. More experimental approaches, such as magnetic and electrically conductive labels, have also surfaced but have yet to mature. Tables 1, 2, and 3 present a summary of assay-based methods developed for micronutrients.
@@
Figure 1: (a) A smartphone camera and LED flash can be used to analyze results from a colorimetric assay. Used with permission from [91]. (b) The design of a colorimetric sweat sensor using microfluidics. Used with permission from [85]. (c) The design of a multiplexed sweat sensor using microfluidics. Used with permission from [86].
3.2.1 Sweat-Based Colorimetric Sensors
@@
Colorimetrics uses reagents to react with an analyte and change color to indicate concentrations of target substances. This color can be analyzed by a camera, such as in a smartphone as demonstrated in Figure 1(a) by Kong et al. [91]. Most sweat-based sensors use microfluidic devices for the collection of sweat [76]. These devices use very tiny valves and channels to "capture and store sweat on the surface of the subject’s skin" via the natural pressure of sweat glands [76]. Sekine et al. [149] demonstrate a microfluidic skin patch with fluorometric probes that target chloride, sodium, and zinc in sweat. A smartphone attachment was designed to take light from the camera’s flash and pass it through an excitation filter, allowing only a particular wavelength through. This excites the fluorometric probes, which emit light that can be captured by the smartphone camera. An emission filter is used to only let fluorescence emission wavelengths through. Similar to other quantitative assays, normalized intensity can be calculated against a reference and used to determine target biomarker concentration. Calibration curves for each nutrient were determined by known concentrations in spiked synthetic sweat. The authors claim strong correlation and accuracy, but do not report exact results. Field studies were conducted where sweat was induced in human volunteers. Measured concentrations were compared to ion chromatography (chloride), atomic absorption spectrometry (AAS; sodium), and inductively coupled plasma mass spectrometry (ICP-MS; zinc). Again, no statistical analysis of results is presented and it appears zinc measurement was the most inaccurate and subject to the most variance.
@@
A paper by Kim et al. [85] describes on-body colorimetric measurement of vitamin C, calcium, zinc, and iron using sweat as a biofluid (Fig. 1(b)). A bespoke colorimetric assay is used for each micronutrient, assessed with known concentrations in a buffer solution. Since temperature and pH can affect colorimetric results, the authors conducted tests within the normal range of body temperature and pH in sweat, finding only a slight shift in results. Uniquely, micronutrients can be supplemented transdermally through the patch itself. Multiple on-body tests were conducted. Sweat was induced by a sauna before and after supplementation (either orally or transdermally). Patch-based measurements were found to be correlated with ICP-MS results. Although the paper claims that "sweat chemistry correlates, at least semiquantitatively, to plasma chemistry" for these micronutrients [85], this claim is based on the time dynamics of concentrations after supplementation rather than a comparison to a gold-standard status assessment (Tables 17 and 19). Furthermore, the clinical literature points out that that vitamin C and iron concentrations in sweat have been shown to have little to no correlation with levels in blood [14, 13]. Therefore, this device provides insights into the rate of excretion of these micronutrients rather than their status. There is no analysis of possible measurement bias induced by transdermal supplementation at the point of sweat collection.
3.2.2 Sweat-Based Multiplexed Sensors

### v2

@@ 3. Biofluid Analysis Methods for Micronutrient Status Assessment
This section describes how biofluid analysis has been leveraged to assess micronutrient status. These assessment methods target particular biomarkers within a biofluid that are indicators of micronutrient status. Reliable biomarkers are a research challenge themselves (which AI techniques may address (Côté and Lamarche, 2022)), but clinical literature suggests that micronutrient biomarkers are more established and specific than macronutrient biomarkers (Elmadfa and Meyer, 2014). Most micronutrients have one or two specific biomarkers associated with their circulating status that are considered to be the optimal reference standard for status assessment (Table LABEL:tab:_assessment). Although some optimal reference standards are still debated, their existence makes the evaluation of novel assessment methods more straightforward.
In this section we will discuss clinical biochemical analysis, followed by other innovative technologies. When reviewing the non-clinical methods (Sections 3.2 to 3.5), we particularly note what each approach claims to assess, the method by which they conduct this assessment, how their method was evaluated, and the clinical relevance of their implementation. Critical to accessibility is each method’s platform. Point-of-care (PoC) devices are compact and portable enough to be deployed for use where needed, as opposed to traditional ‘benchtop’ technologies that are restricted to a laboratory setting. The term PoC is widely applicable, so we label methods more specifically to compare the advantages and disadvantages of each, while still acknowledging that they are considered PoC. Specifically, the terms ‘portable’, ‘smartphone-based’, and ‘wearable’ all imply PoC devices but also suggest varying levels of accessibility and ubiquity.
@@
We further expand on clinical relevance by noting that this includes the target biomarker, the assessment method, and the concentrations of that biomarker that are evaluated. A clinically relevant method should closely align with the optimal reference standard on these factors (Table LABEL:tab:_assessment). To aid future work that may wish to integrate or innovate on a particular method, we explicitly mention when a study does not demonstrate this agreement (e.g. assessing RBP for vitamin A status) and/or what assessment methods were used during evaluation (e.g. ELISA). For micronutrients with rapid turnover (i.e. water-soluble vitamins) or without risks of excess, a lack of sensitivity in the upper spectrum of the clinically relevant concentrations (Table LABEL:tab:_assessment) are not a major limitation for methods which primarily aim to identify deficiency. Regardless, such a limitation is still noted for completeness. Last, we provide tables that group together similar works and summarize the pertinent details and quantitative results of their evaluation, if available. For conciseness, this information is omitted from the body of the text, and we encourage readers to instead refer to the relevant tables.
3.1. Clinical Biochemical Analysis
Clinical biochemical analysis involves laboratory testing of biomarkers found in urine, blood, or other biosamples (Reber et al., 2019). Results can be influenced by several factors and need to be interpreted in the context of other aspects of the patient’s health. Additionally, biochemical testing is often time- and resource-intensive. Despite this, biochemical analysis describes clinical optimal reference standard methods for quantifying the circulating micronutrient status in the body (Zhang et al., 2018; Höller et al., 2018). These methods can be roughly separated into various types of assays, and liquid chromatography (LC)-coupled spectroscopy.
One type of assay which is popular for clinical micronutrient assessment is a microbiological inhibition assay. These assays work on the principle that specific micronutrients are needed for the growth of certain bacteria, and this growth can be measured to indicate the amount of a micronutrient present in a sample (Zhang et al., 2018). Microbiological inhibition assays were previously the widely-accepted optimal reference standard, but improvements in LC and spectroscopy highlight their relatively poor precision and accuracy. These flaws have relegated microbiological inhibition assay methods to be used mostly during screening or in resource-constrained testing, except for some micronutrients, where they remain the standard. Another common type of assay applied in studies is the enzyme-linked immunosorbent assay (ELISA). An ELISA test is used for measuring antibodies in blood, and is a useful clinical screening tool for further testing (Kinman, 2012). While not the optimal reference standard for the assessment of micronutrient status, ELISA tests can yield valuable data for studies that are time or resource limited. Lastly, we find that antibody or immunoassays as well as colorimetric and fluorometric assays are popular in emerging accessible and non-invasive assessment technologies.
Modern optimal reference standards overwhelmingly apply LC-coupled spectroscopy (Zhang et al., 2018; Höller et al., 2018) (Table LABEL:tab:_assessment). LC is defined as ”a separation process used to isolate the individual components in a mixture” (Chemyx, 2017). High-performance LC (HPLC) uses pressure to facilitate the separation process, reducing the time required. It is commonly coupled with mass spectroscopy (MS) in optimal reference standard approaches (Zhang et al., 2018; Höller et al., 2018). Spectroscopy is the ”investigation and measurement of spectra produced by matter interacting with or emitting electromagnetic radiation” (noa, 2019). Every molecule reacts to the applied radiation in a unique way that allows us to ”detect, determine, or quantify the molecular and/or structural composition of a sample” (noa, 2019). MS is the most important subfield of spectroscopy to understand for biochemical analysis. It measures the mass-to-charge ratio of the molecules in a sample as a way to determine and quantify the composition of molecules in the sample (noa, 2019). This is done by vaporizing the molecules in a sample into gas-phase ions, which are then sorted by their mass-to-charge ratios. We will discuss other forms of spectroscopy and their utility for micronutrient status assessment further in Section 3.4.
Matrices are the biosamples that are the subjects of the aforementioned methods of analysis. Most are blood based, but in a few cases urine is used in the optimal reference standard (mostly for water-soluble vitamins) (Höller et al., 2018). Blood matrices are whole blood, washed red blood cells, plasma, and serum. Whole blood is blood as it is from the vein (venous blood). Washed red blood cells are red blood cells that have been separated from the other components of blood such as plasma, platelets, and white blood cells (Keir et al., 2016). Plasma is obtained by adding an anticoagulant to whole blood and placing it in a centrifuge (Systems, 2021). Serum is obtained similarly to plasma, except the blood is allowed to clot before centrifuging. However, blood can also be obtained from the capillaries (capillary blood) as opposed to the vein, often via a finger prick. Capillary blood is often easier and cheaper to obtain, especially via untrained personnel, but contains a mix of venous and arterial blood, together with interstitial fluid which surrounds cells in the body (Royal et al., 2022). Because of this, test results on capillary vs venous blood can differ (e.g. hemoglobin concentrations are higher in capillary blood) and so the two should not be considered interchangeably.
Table 1. Optimal Reference Standard Methods of Assessing Micronutrient Imbalance. Information from (Höller et al., 2018; Zhang et al., 2018; Mueller, 2017; Berger et al., 2022; National Institutes of Health, datea).
Micronutrient Method Biomarker Matrix Intervals Impact of Inflammation Approx. Cost (US; Walk-In Lab)
Vitamin B1 Erythrocyte transketolase activity coefficient assay Increase in erythrocyte transketolase activity Washed red blood cells Deficient: ¿25%; Insufficient: 15-25%; Sufficient: ¡15% None on direct plasma levels \$65
Vitamin B2 Erythrocyte glutathione reductase activity coefficient assay Increase in erythrocyte glutathione reductase activity Washed red blood cells Deficient: ¿40%; Insufficient: 20-40%; Sufficient: ¡20% Decrease in plasma levels (erythrocyte assays are more stable) \$120
Vitamin B3 LC-MS/MS Niacin metabolites (NMN and 2-pyr, limited representation of stores and recent intake) Urine Deficient: ¡5.8; Insufficient: 17.5-5.8; Sufficient: ¿17.5 mcmol/day Lack of evidence \$179
Vitamin B5 LC-MS/MS Pantothenic acid (requires enzyme pretreatment) Whole blood Deficient: ¡0.22; Sufficient: 0.35-0.59 mg/L Lack of evidence \$149
Vitamin B6 HPLC or LC-MS/MS pyridoxal phosphate (PLP) Plasma or serum Sufficient: ¿4.94 or ¿7.41 $\mu$g/L plasma PLP (varies by source) Decrease in plasma PLP, no effect on RBC concentration \$69
Vitamin B7 (Eng et al., 2013) LC-MS/MS for biotin; gel densitometry for MCC/PCC (Mall et al., 2010) Biotin (less sensitive); holo-MCC and holo-PCC (only reliable markers) Urine for biotin; WBCs for MCC/PCC Sufficient: 4.4-31 $\mu$g/day urinary biotin, 8.2 arbitrary units holo-MCC, 9.1 arbitrary units holo-PCC None on biomarkers \$199
Vitamin B9 LC-MS/MS Folate Serum for altered exposure and recent intake; Red blood cell for long term/3 month status and storage levels Sufficient: ¿3 ng/mL serum, ¿140 ng/mL RBC Lack of evidence \$29
Vitamin B12 (Ada’s Medical Knowledge Team, 2023) GC-MS B12, confirmed with methylmalonic acid (MMA, also related to B2, B6, folate); no single ’optimal reference standard’ Plasma or serum Deficient: ¡200-250 pg/mL B12, ¿0.03 mg/L MMA (some debate over this) Association with increased B12 levels \$35
Vitamin C HPLC Ascorbate Plasma (some claim serum should be avoided) Deficient: 1.94; Insufficient: 2.11-4.05; Sufficient: 4.05 mg/L Decrease in plasma ascorbic acid (rapid, decrease when CRP ¿10 mg/L, normal values not detected if CRP ¿40 mg/L) \$49
Vitamin A LC-MS/MS Retinol (only sensitive to deficiency or excess in storage, affected by infection and protein/zinc deficiency). Best method is to indirectly measure reserves in liver over several days of administration Plasma or serum Severely deficient: ¡0.1; Deficient: 0.1-0.2; Sufficient: 0.3-1; Toxic: ¿1 mg/L retinol Decrease in serum retinol (adjustment equations exist but are not universally applicable, e.g. BRINDA (Luo et al., 2016)) \$58
Vitamin D (Ross and Institute of Medicine (2011), U. S.) LC-MS/MS 25(OH)D (calcidiol) Plasma or serum Deficient: ¡12; Insufficient: 12-¡20; Sufficient: 20-50; Toxic: ¿50 ng/mL (not definitively established/linked to clinical outcomes, varies based on assay and lab) Decrease in plasma levels (all values below reference ranges with CBP ¿40 mg/L) \$59
Vitamin E LC-UV Ratio of Vit E to total blood lipids Plasma or serum Insufficient: ¡5.17 mg/L Vit E, ¡0.8 mg Vit E/g total lipid; Sufficient: 8.6-13 mg/L Vit E (adults have higher levels) Some effects (blood concentrations less interpretable at CRP ¿80 mg/L) \$48
Vitamin K Immuno-based assays Plasma phylloquinone (usually for short term intake, no single ‘optimal reference standard’); prothrombin time (time to blood clot, only clinically relevant measure); variety of other ’functional’ biomarkers Plasma Deficient: ¡0.15; Sufficient (fasting): 0.15-1 $\mu$g/L Status associated with lower inflammatory marker concentration \$96
Iron (Gloria, date; Gerber, 2024) Electrochemi-luminescence immunoassay (ECLIA) Ferritin for deficiency (first phase, evaluates storage, inflated by infection); Iron increase after supplementation for malabsorption; Hemoglobin used to confirm IDA Serum Iron-deficiency anemia: ¡10, Deficiency: 10-30 $\mu$g/L ferritin Ferritin may be inflated, falsely normal/misleading (adjustment equations exist but are not universally applicable, e.g. BRINDA (Luo et al., 2016)) \$29
Copper ICP-MS Copper or ceruloplasmin (CP), neither reliable Serum Depletion: ¡50.8 (copper); Deficient: 50.8-76.2 (copper, high CRP); Sufficient: 63.5-158.9 $\mu$g/dL (copper), 180-400 mg/L (CP) Increase in plasma concentrations \$33
Zinc (Pat Bass, 2023) Atomic Absorption Spectroscopy (AAS) Zinc (levels are halved by Systemic Inflammatory Response Syndrome, can be normal with clinical symptoms present, levels vary with time of day so it is recommended that albumin and CRP changes are taken into account) Plasma or serum Deficient: 70 women, 74 men; Insufficient: 70/74-80; Sufficient: 80-120 $\mu$g/dL Decrease in plasma levels (significant when CRP exceeds 20 mg/L, adjustment equations exist but are not universally applicable, e.g. BRINDA (Luo et al., 2016)) \$38
Iodine ICP-MS Iodine Urine (24h or random), serum less recommended Depletion: ¡20, NA; Deficient: 20-100, ¡40; Sufficient: 100-300 $\mu$g/24hr urine, 40-100 $\mu$g/L serum (levels should be higher in those who are pregnant or lactating) Lack of evidence \$89
Selenium AAS, ICP-MS Selenium (recent intake) or Selenoprotein P Plasma or serum Deficient: ¡60; Sufficient: ¿60; Toxicity: ¿474 to 948 $\mu$g/L Se (intervals vary by source and population: women and black people have naturally lower concentrations) Decrease in plasma levels proportional to inflammation, can be adjusted for \$99
Magnesium AAS Magnesium Serum (little correlation with overall status or tissue stores) and urine (after supplementation) Deficient: ¡18.23; Sufficient: 18.23-23.1 mg/L serum Mg \$28
3.2. Assay-Based Technology
@@
One of the largest areas of work applicable to micronutrient status explored by this review is quantitative assays. Although we cover many different types of quantitative assays, a comprehensive review of lateral flow quantitative assays by Urusov et al. (2019) provides a robust background for how these devices work. Lateral flow (immunochromatographic) assays indicate that a target compound is either present in the sample or present in excess of a particular threshold, usually via staining on the test membrane. These types of assays are useful for tests that benefit from quick conclusions (e.g. pregnancy tests). The most prevalent approach for extracting quantitative information from lateral flow assays is via optical signal registration (Urusov et al., 2019). This method involves the analysis of absorbed and reflected light from the test surface and the staining upon it, similar to the practice of spectrophotometry. The test/control ratio is a common metric used to quantify the magnitude of this staining relative to a control or reference area. Some commercially-available devices are limited to automatically confirming the presence of the test line, while others can use line intensity to calculate analyte content. Urusov et al. (2019) note that portability has become a recent focus in this market, and this is not just limited to specialized devices in a portable platform. Smartphones have been successfully used for optical signal registration, even with fluorescent labels that decrease detection limits. Some manufacturers provide their own smartphone apps for quantitative analysis, but controlling for lighting and positioning is a challenge. To address this issue, another approach is the use of a standardized scanning device to collect image data and the off-device analysis of the image by specialized software. More experimental approaches, such as magnetic and electrically conductive labels, have also surfaced but have yet to mature. Tables 2, 3, and 4 present a summary of assay-based methods developed for micronutrients.
@@
Figure 1. (i) A smartphone camera and LED flash can be used to analyze results from a colorimetric assay. Used with permission from Kong et al. (2019). (ii) The design of a multiplexed sweat sensor using microfluidics. Used with permission from Kim et al. (2020). (iii) The design of a colorimetric sweat sensor using microfluidics. Used with permission from Kim et al. (2022).
[…]
(i)(a) Graphic describing the process of taking a sample, adding it to the test area of the assay, capturing an image of the assay’s response to the sample with a smartphone, then analyzing the image using the test spots and black/white reference areas embedded into the test. (b) A smartphone screen is shown taking an image of the test area with the flash on. Beside, a heatmap of light intensity from the captured image. (c) A comparison of images of the test area, taken with Huawei, iPhone, and Samsung brand smartphones, with the phone’s LED flash one or off, under two external lighting conditions. (ii)(a) Exploded-view of the skin-interfaced, multiplexed assay showing its electronic componentry. (b) Labeled diagram of the skeleton microfluidics and electrodes. The sample enters through an inlet, travels through the microfluidic channels to interact with the electrodes, then exits out a ventilation port. (c) Labeled diagram showing the tolerances of the skeleton microfluidics and electrodes. (d) Labeled diagram of the device’s integrated LFIA strip, showing the capillary valves, sampling reservoir, and reference area. (e) Three images of the device being handled and bent. (f) Corresponding strain maps of the device being handled and bent. (iii)(a) Exploded-view of a skin-interfaced colorimetric sweat sensor, highlighting the channels, chemical reagents, reference areas, and nutrient supplementation patches. (b) Labeled diagram showing the shape and size of a microfluidic reservoir. (c) The sweat sensor is shown fitting easily on someone’s fingertip. (d) The sweat sensor held in someone’s palm, next to a pill for scale. (e) The nutrient patch is pulled back to reveal the block layer between the assays and the supplementation patch. For more information on each figure, please refer to their respective sources.
3.2.1. Sweat-Based Colorimetric Sensors
@@
Colorimetrics uses reagents to react with an analyte and change color to indicate concentrations of target substances. This color can be analyzed by a camera, such as in a smartphone as demonstrated in Figure 1(i) by Kong et al. (2019). Most sweat-based sensors use microfluidic devices for the collection of sweat (Jo et al., 2021). These devices use very tiny valves and channels to ”capture and store sweat on the surface of the subject’s skin” via the natural pressure of sweat glands (Jo et al., 2021). Sekine et al. (2018) demonstrated a microfluidic skin patch with fluorometric probes that aimed to analyze chloride, sodium, and zinc in sweat. A smartphone attachment was designed to take light from the camera’s flash and pass it through an excitation filter, allowing only a particular wavelength through. Similar to other quantitative assays, normalized intensity can be calculated against a reference and used to determine biomarker concentration and calibration curves for each nutrient were determined by known concentrations in spiked synthetic sweat. The authors claimed a strong correlation and accuracy, but do not report exact results. Field studies were conducted where sweat was induced in human volunteers (undisclosed sample size). Measured concentrations were compared to ion chromatography (chloride), atomic absorption spectrometry (AAS; sodium), and inductively coupled plasma mass spectrometry (ICP-MS; zinc). Again, no statistical analysis of results is presented and it appears zinc measurement was the most inaccurate and subject to the most variance.
@@
A paper by Kim et al. (2022) described on-body colorimetric measurement of vitamin C, calcium, zinc, and iron using sweat as a biofluid (Fig. 1(ii)). A bespoke colorimetric assay was used for each micronutrient, assessed with known concentrations in a buffer solution. Since temperature and pH can affect colorimetric results, the authors conducted tests within the normal range of body temperature and pH in sweat, finding only a slight shift in results. Uniquely, micronutrients could be supplemented transdermally through the patch itself. Multiple on-body tests were conducted with 7 volunteers (4 male). Sweat was induced by a sauna before and after supplementation (either orally or transdermally). Patch-based measurements were found to be correlated with ICP-MS results. Although the paper claimed that ”sweat chemistry correlates, at least semiquantitatively, to plasma chemistry” for these micronutrients (Kim et al., 2022), this claim is based on the time dynamics of concentrations after supplementation rather than a comparison to an optimal reference standard status assessment (Table LABEL:tab:_assessment). Furthermore, the clinical literature points out that that vitamin C and iron concentrations in sweat have been shown to have little to no correlation with levels in blood (Baker and Wolfe, 2020; Baker, 2019). Therefore, we argue that this device provides insights into the rate of excretion of these micronutrients rather than their status. Lastly, there is no analysis of possible measurement bias induced by transdermal supplementation at the point of sweat collection.
3.2.2. Sweat-Based Multiplexed Sensors
Multiplexed analyses combine and analyze data from assays with multiple sensors such as electrocardiogram (ECG), temperature, electrodermal activity (EDA), HRV data and more. Thus, these multiplexed sensors can enable more complex physiologic monitoring and diagnosis. For example, EDA measures the change in skin conductance caused by sweat, an indicator of nervous system arousal (Posada-Quintero and Chon, 2020), which can be associated with micronutrient status (Section 2.3).